### v2

# Never-Ending Behavior-Cloning Agent for Robotic Manipulation

Abstract
Relying on multi-modal observations, embodied robots could perform multiple robotic manipulation tasks in unstructured real-world environments. However, most language-conditioned behavior-cloning agents still face existing long-standing challenges, i.e., 3D scene representation and human-level task learning, when adapting into new sequential tasks in practical scenarios. We here investigate these above challenges with NBAgent in embodied robots, a pioneering language-conditioned Never-ending Behavior-cloning Agent. It can continually learn observation knowledge of novel 3D scene semantics and robot manipulation skills from skill-shared and skill-specific attributes, respectively. Specifically, we propose a skill-shared semantic rendering module and a skill-shared representation distillation module to effectively learn 3D scene semantics from skill-shared attribute, further tackling 3D scene representation overlooking. Meanwhile, we establish a skill-specific evolving planner to perform manipulation knowledge decoupling, which can continually embed novel skill-specific knowledge like human from latent and low-rank space. Finally, we design a never-ending embodied robot manipulation benchmark, and expensive experiments demonstrate the significant performance of our method.
*Corresponding Author.
1 Introduction
Robot learning has attracted growing interests in integrating machine learning with robot control system to solve various robotic tasks, such as manipulation [39], navigation [36], mapping and localization [27]. In recent decades, behavior-cloning agents [17, 53] have demonstrated notable success in effective training with only a few demonstrations and direct implementation on real robots. Benefiting from vision-language observations, current multi-task behavior-cloning methods focus on utilizing multi-modal data to efficiently execute complex manipulation tasks with visual observations. For instance, PerAct [39] utilizes a PerceiverIO Transformer [21] to encode language goals and RGB-D voxel observations, subsequently generating discretized robotic actions.
However, most existing language-conditioned behavior-cloning methods [21, 17, 26] assume that the skills mastered by a robot remain unchanged over time in real-world 2D scenarios, and focus on training policy model on a fixed set of manipulation tasks. They are resource-constrained when undertaking new skills and handling novel objects with intricate structures [44]. To address this scenario, the embodied robots—like human—should be capable of learning novel challenging manipulation skills in a continual learning manner. For instance, a home robot in the open-ended world is expected to consecutively learn various novel manipulation skills with demonstrations of behavior, and timely meet the evolving needs of its owners. Generally, a trivial approach for this scenario involves retraining the robot on all past and novel data, which leads to a large computational burden and high cost of memory storage, thereby is limited in real world unstructured 3D applications.
To address the aforementioned scenarios, we consider a practical challenging robot learning problem, i.e., Never-ending Behavior-cloning Robot Learning (NBRL), where an agent can perform continual learning on successive behavior-cloning manipulation tasks and counteract catastrophic forgetting from learned skills. Different from learning category-wise knowledge focused by most existing methods [33, 44, 11], we here take the attempt to rethink some attributes about skill-wise knowledge:
•  Skill-Shared Attribute indicates that different complex skills possess shared knowledge, such as similar manipulated objects and 3D scenes understanding and so on. We primarily investigate the consistency of skill-shared knowledge across various skills on semantics of 3D scenes, which plays a key role in addressing skill forgetting and achieving comprehensive understanding on 3D scenes.
•  Skill-Specific Attribute originates from various factors such as distinct manipulation sequences, object recognition, motion primitives, and other distinct elements within each unique skill. Focusing the embodied robot to learn skill-specific knowledge can effectively perform novel skills learning and tackle forgetting on past learned manipulation skills or tasks.
To tackle the above-mentioned challenges, we propose a pioneering language-conditioned Never-ending Behavior-cloning Agent (i.e., NBAgent), which can continually acquire skill-wise knowledge from both skill-shared and skill-specific attributes. To the best of our knowledge, this is an earlier attempt to explore continual learning for multi-modal behavior-cloning robotic manipulation. To be specific, we design a skill-shared semantic rendering module (SSR) and a skill-shared representation distillation module (SRD) to transfer skill-shared knowledge on semantics of 3D scenes. Supervised by Neural Radiance Fields (NeRFs) and a vision foundation model, SSR can transfer skill-shared semantic from 2D space into 3D space across novel and old skills; SRD can effectively distill skill-shared knowledge between old and current models to well align voxel representation. Additionally, we propose a skill-specific evolving planner (SEP) to decouple the skill-wise knowledge in the latent and low-rank space, and focus on skill-specific knowledge learning to alleviate manipulation skill forgetting. Several major contributions of our work are as follows:
•  We take the earlier attempt to explore a practical challenging problem called Never-ending Behavior-cloning Robot Learning (NBRL), where we propose Never-ending Behavior-cloning Agent (i.e., NBAgent) to address the core challenges of skill-wise knowledge learning in 3D scene from skill-shared and skill-specific attributes.
•  We develop a skill-shared semantic rendering module and a skill-shared representation distillation module to capture the 3D semantic knowledge, which can transfer skill-shared semantic of 3D scenes to overcome 3D reasoning overlooking in continual learning.
•  We design a skill-specific evolving planner to perform skill-specific knowledge learning, which can decouple the skill-wise knowledge into latent and low-rank space, and continually embed novel skill-specific knowledge. Comprehensive experiments conducted on our proposed NBRL benchmark for home robotic manipulation demonstrate the effectiveness and robustness of our method.
2 Related Work
Robotic Manipulation. Recent works [6, 8, 20, 37, 14, 5, 54] have resulted in substantial advancements in the accomplishment of intricate tasks. VIMA [24] proposes a novel multi-modal prompting scheme, which transforms a wide range of robotic manipulation tasks into a sequence modeling problem. Compared with directly using images as the manipulation input [16, 38], voxelizing 3D point clouds as a 3D representation [39, 53, 17, 23] can accomplish more complex tasks. PerAct [39] enables agent to perform better in robotic manipulation by voxelizing RGB-D images and discretizing output actions. GNFactor [53] can generalize neural radiance field rendering through joint training, improving the model’s generalization ability.
Continual Learning. Continual learning provides the foundation for the adaptive development of AI systems [46, 10]. The main approaches of continual learning can be categorized into three directions: Parameter regularization-based methods [33, 30, 9, 12, 11] balance the old and new tasks by adding more explicit regularization terms. Architecture-based methods [25, 48, 45, 43] construct network parameters for different tasks. Replay-based methods include empirical replays [4, 33, 41, 42] and generative replays [29, 49, 40]. Some works focus on the improvement of robots by continual learning [3, 15, 18, 2, 1]. LOTUS [44] stores a few human demos of novel tasks into the growing skill library, enabling lifelong learning ability for robots. However, these methods are build on 2D vision observations, and cannot be applied to perform language-conditioned behaviour-cloning manipulation with multi-modal data.
3 Methodology
3.1 Problem Definition and Overview
Problem Definition. By following traditional continual learning methods [33, 12], we define a multi-modal manipulation skill data stream as $\mathcal{T}=\{\mathcal{T}^{m}\}_{m=1}^{M}$, where $M$ denotes the number of continual skill learning tasks. The objective of NBAgent is to execute all acquired skills after observing $\mathcal{T}$. Specifically, each skill learning task consists of various robotic manipulation skills. The $m$-th continual skill learning task $\mathcal{T}^{m}=\{\mathcal{D}^{m}_{i}\}_{i=1}^{N^{d}}$ consists of $N^{d}$ skill demonstrations. Each skill demonstration $\mathcal{D}^{m}_{i}$ can be extracted to a set of keyframe actions [39], i.e., $\mathcal{D}^{m}_{i}=\{\mathbf{k}^{m,i}_{j}\}_{j=1}^{N^{k}},\mathbf{k}^{m,i}_{j}=\{\mathbf{a}^{m,i}_{j},\mathbf{r}^{m,i}_{j},\mathbf{l}^{m,i}\}$, where $N^{k}$ is the total keyframe action quantity. Given the current state in action space $\mathbf{a}$, the structured observation $\mathbf{r}$ and language instruction $\mathbf{l}$, the agent is expected to predict the next best keyframe action, which can be served as an action classification task [23]. Additionally, the structured observation $\mathbf{r}$ is composed of the RGB-D images captured by a single front camera. An action state $\mathbf{a}$ can be divided into a discretized translation $\mathbf{a}_{tran}\in\mathbb{R}^{3}$, rotation $\mathbf{a}_{rot}\in\mathbb{R}^{(360/5)\times 3}$, gripper open state $\mathbf{a}_{grip}\in\{0,1\}$, and collision avoidance $\mathbf{a}_{col}\in\{0,1\}$. Here the rotation parameter $\mathbf{a}_{rot}$ entails the discretization of each rotation axis into a set of R = 5 bins. The collision avoidance parameter $\mathbf{a}_{col}$ provides guidance to the agent regarding the imperative need to avoid collisions.
Figure 1: Overview of the proposed NBAgent. It consists of a skill-shared semantic rendering module and a skill-shared representation distillation loss $\mathcal{L}_{\mathrm{SRD}}$ to transfer skill-shared knowledge on semantics of 3D scenes, and a skill-specific evolving planner to learn skill-specific knowledge.
Overview. The overview of NBAgent to learn skill-wise knowledge is shown in Fig. 1. When observing a novel sequential skill learning task $\mathcal{T}^{m}$, we initialize the perceiver model $\Theta^{m}_{p}$ for the current task utilizing the model $\Theta^{m-1}_{p}$ obtained from the last task, and store $\Theta^{m-1}_{p}$ as a teacher model to compute the SRD loss. Given a RGB-D image and language instruction, as shown in Fig. 1, $\Theta^{m}_{p}$ firstly encodes RGB-D input to obtain a deep 3D voxel via utilizing a scaled-down voxel encoder. We here design a SSR module to transfer skill-shared semantic of 3D scenes across novel and past encountered skills. Afterwards, the patched voxel and language embeddings are input to cross-attention blocks and self-attention blocks to perform feature extraction and semantics fusion, where we propose SEP to learn skill-specific knowledge from latent and low-rank space. Finally, we utilize a Q-function head $\mathbf{Q}^{m}$ to predict the state of the next keyframe in voxel space, where a SRD loss $\mathcal{L}_{\mathrm{SRD}}$ is developed to tackle catastrophic forgetting by aligning skill-shared voxel representation.
3.2 Skill-Shared Semantic Rendering Module
For language-conditional behaviour-cloning manipulation, a comprehensive semantics understanding of the 3D scenes [13] plays a key role in enabling agent to perform complicated manipulation skills. Especially in NBRL problem, which exists skill-shared semantic across various skills, such as 3D objects and scene semantics. The existence of overlooking on common semantic space makes these semantics incomplete, further resulting in catastrophic forgetting on past learned skills. Considering this motivation, we develop a skill-shared semantic rendering module (SSR) to transfer skill-shared semantic of 3D scenes, where a NeRF model and a vision foundation model are introduced to provide semantics supervision for this transfer process.
Concretely, we draw inspiration from 3D visual representation learning [53], and leverage a latent-conditioned NeRF architecture [52]. This architecture could synthesize RGB color $\mathbf{c}$ of a novel image views like traditional NeRF [31], and render the semantics $\mathbf{s}$ from the 3D voxel space as follows:
$\mathcal{F}_{\Theta^{m}_{n}}(\mathbf{x},\mathbf{d},\mathbf{v}_{s})=(\sigma,\mathbf{c},\mathbf{s}),$ (1)
[…]
To distill skill-shared semantic in the $m$-th skill learning task, we initialize a NeRF model acquired from the last task and denote it as $\Theta_{n}^{m-1}$. Then, we feed the same input to obtain the pseudo ground truth $\hat{\mathbf{C}}$ by Eq. (2). We design a loss function to supervise the reconstruction process as follows:
$\mathcal{L}_{\mathrm{C}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\mathbf{Y}_{c}(\mathbf{r})\|^{2}_{2}+\beta\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{C}}(\mathbf{r},\mathbf{v}_{s})\|^{2}_{2}\cdot\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}},$ (3)
where $\mathbf{Y}_{c}$ indicates the ground truth color and $\mathcal{R}$ is the set of all camera rays. $\alpha$ is the hyper-parameter to control the weight of loss function. $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}$ is defined such that $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}=1$, when the condition $\mathbf{v}_{s}\notin\mathcal{T}^{m}$ is satisfied, and $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}=0$ otherwise.
Considering the insufficiency in capturing skill-shared semantic by reconstructing novel views, we introduce a pre-trained visual foundation model that contains robust scene semantics to provide supervision. Relying on being pre-trained on large-scale vision-language dataset, Stable Diffusion model [35] can possess robust intrinsic representational capabilities, which is consequently utilized for semantic representation in segmentation and classification tasks [50, 28]. In light of this, we employ Stable Diffusion model $\Theta_{u}$ to extract vision-language semantics for supervision. Given a input view, i.e., $\mathbf{Y}_{c}$, we perform a one-step noise adding process to obtain a noisy image $\mathbf{Y}_{c,t}$. Then we utilize diffusion model $\Theta_{u}$ to collect vision-language semantic feature as the ground truth $\hat{\mathbf{F}}_{s}$:
$\mathbf{Y}_{c,t}(\mathbf{r}):=\sqrt{\alpha_{t}}\mathcal{E}_{v}(\mathbf{Y}_{c}(\mathbf{r}))+\sqrt{1-\alpha_{t}}\epsilon,~{}~{}\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})=\Theta_{u}(\mathbf{Y}_{c,t}(\mathbf{r}),\mathcal{E}_{c}(\mathbf{l}_{p})),$ (4)
where $\mathcal{E}_{v}$ is a VAE encoder to encode image $\mathbf{Y}_{c}$ from pixel space to latent semantic space. $t$ represents the diffusion process step, $\epsilon\sim\mathcal{N}(0,1)$ and $\alpha_{t}$ is designed to control the noise schedule. $\mathbf{l}_{p}$ denotes the language prompt modified from task description $\mathbf{l}$. To perform skill-shared semantic transfer, we align the rendered semantic feature $\mathbf{F}_{s}$ and diffusion feature $\hat{\mathbf{F}}_{s}$, and obtain the major objective of our SSR module as follows:
$\mathcal{L}_{\mathrm{SSR}}=\mathcal{L}_{\mathrm{C}}+\lambda_{1}\mathcal{L}_{\mathrm{S}},~{}~{}\mathcal{L}_{\mathrm{S}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{F}_{s}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})\|^{2}_{2}.$ (5)
Algorithm 1 Optimization Pipeline of NBAgent.
0:  Continual skill learning tasks $\{\mathcal{T}^{m}\}_{m=1}^{M}$ with datasets $\mathcal{T}^{m}=\{\mathcal{D}^{m}\}_{i=1}^{N^{d}}$; Initialized perceiver model: $\Theta^{0}_{p}$; Initialized NeRF model: $\Theta^{0}_{n}$; Pre-trained diffusion model: $\Theta_{u}$; Pre-trained CLIP language encoder: $\mathcal{E}_{c}$; Initialized memory buff: $\mathcal{M}=\emptyset$; Iterations: $\{\mathcal{I}^{m}\}_{m=1}^{M}$.
1:  #While observing a new task $\mathcal{T}^{m}$:
2:  for $z=1,2,\cdots,\mathcal{I}^{m}$ do
3:     Randomly select keyframe $\mathbf{k}^{m}_{j}$ from $\{\mathcal{D}_{i}^{m}\}_{i=1}^{N^{d}}\cup\mathcal{M}$;
4:     Obtain $\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{s},\mathrm{l}_{x}$ utilizing $\mathcal{E}^{m}_{v}$, $\mathcal{E}^{m-1}_{v}$and $\mathcal{E}_{c}$;
5:     Compute $\mathcal{L}_{\mathrm{SSR}}$ by SSR ($\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{x},\Theta_{u}$) using Eq. (1);
6:     $\mathbf{S}[h,:],\mathbf{W}[h,:]\xleftarrow{}$ SEP ($\mathbf{l}_{s}$);
7:     Compute $\mathcal{L}_{\mathrm{CE}},\mathcal{L}_{\mathrm{SRD}}$ utilizing $\mathbf{S}[h,:]$, $\mathbf{W}[h,:]$, $\Theta^{m}_{p},\Theta^{m-1}_{p}$;
8:     Update $\Theta^{m}_{p}$ by Eq. (11);
9:  end for
10:  Store few samples from $\{\mathcal{D}_{i}^{m}\}_{i=1}^{N^{d}}$ in $\mathcal{M}$;
11:  Return: $\Theta^{m}_{p},\mathcal{M}$.
3.3 Skill-Shared Representation Distillation Module
To address semantic overlooking on past learned skills, we develop a skill-shared representation distillation module (SRD) to align skill-shared semantic in 3D voxel representation space, as presented in Fig. 1. Specifically, given a multi-modal keyframe input $\mathbf{k}^{m}_{j}=\{\mathbf{a}^{m}_{j},\mathbf{r}^{m}_{j},\mathbf{l}^{m}\}$ from a mini-batch, we feed it into our perceiver model $\mathrm{\Theta}^{m}_{p}$ to obtain a keyframe on voxel space, denoted as voxel representation $\mathbf{v}^{m}_{r,j}$. The next keyframe prediction can then be computed utilizing a Q-function head as $\mathbf{Q}^{m}(\mathbf{v}^{m}_{r,j})=\{\mathbf{P}^{m}_{j,tran},\mathbf{P}^{m}_{j,rot},\mathbf{P}^{m}_{j,grip},\mathbf{P}^{m}_{j,col}\}$, where $\mathbf{P}^{m}_{j,tran},\mathbf{P}^{m}_{j,rot},\mathbf{P}^{m}_{j,grip},\mathbf{P}^{m}_{j,col}$ denote prediction on discretized translation, rotation, gripper open state and collision avoidance. To supervise NBAgent to learn skill-wise knowledge, we follow [39], and introduce a cross-entropy loss as :
$\mathcal{L}_{\mathrm{CE}}=-\frac{1}{B}\sum^{B}_{j=1}\mathbf{Y}^{m}_{j}\mathrm{log}(\mathbf{Q}^{m}(\mathbf{v}^{m}_{r,j}),~{}~{}\mathbf{v}^{m}_{r,j}=\Theta_{p}^{m}(k^{m}_{j}),$ (6)
where $B$ represents the batch size, and $\mathbf{Y}^{m}_{j}=\{\mathbf{Y}^{m}_{j,tran},\mathbf{Y}^{m}_{j,rot},\mathbf{Y}^{m}_{j,grip},\mathbf{Y}^{m}_{j,col}\}$ is the ground truth for predicting the next keyframe. In light of this, NBAgent can continually learn skill-wise knowledge from current dataset $\mathcal{T}^{m}$ and memory buff $\mathcal{M}$. However, due to the limited amount of data available from old skills in memory buff $\mathcal{M}$, the skill-shared semantic drift on 3D voxel representation occurs between novel and old skills. It further results in forgetting skill-shared knowledge on old skills.
To address the aforementioned problems, we take an attempt to employ knowledge distillation to align voxel representation between old and new model in NBRL problem. Specifically, we initialize a teacher model with the perceiver model $\Theta^{m-1}_{p}$ from the last task to extract the soft label: $\hat{\mathbf{Y}}^{m}_{j}=\mathbf{Q}^{m-1}(\mathbf{v}^{m-1}_{r,j}),\hat{\mathbf{Y}}^{m}_{j}=\{\hat{\mathbf{Y}}^{m}_{j,tran},\hat{\mathbf{Y}}^{m}_{j,rot},\hat{\mathbf{Y}}^{m}_{j,grip},\hat{\mathbf{Y}}^{m}_{j,col}\}$ and apply the Kullback-Leibler divergence to align the outputs of two agents as follows:
$\mathcal{L}_{\mathrm{SRD}}=\frac{1}{\hat{B}}\sum^{B}_{j=1}\rho(\hat{\mathbf{Y}}^{m}_{j}/\tau)\mathrm{log}(\frac{\rho(\hat{\mathbf{Y}}^{m}_{j}/\tau)}{\rho(\mathbf{Q}^{m}(\mathbf{v}^{m}_{r,j}/\tau)})\cdot\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}},$ (7)
where $\hat{B}=\sum_{j=1}^{B}\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}}$, and $\tau$ is a hyper-parameter represents distillation temperature.
3.4 Skill-Specific Evolving Planner
To learn category-wise knowledge over the above 3D scene representation, existing continual learning methods [33, 12] assume that the knowledge acquired from novel tasks and that from previous tasks are mutually independent. Differently, we consider decoupling the knowledge as the skill-shared knowledge and skill-specific knowledge, when learning skill-wise knowledge. For instance, an agent aims to “stack the wine bottle” after learning how to “open the wine bottle”. It does not need to relearn the skill-shared knowledge of 3D scene understanding and object recognition; instead, the agent is expected to focus on acquiring the skill-specific knowledge related to the operation sequence of stacking and reducing forgetting on skill-shared knowledge. Considering this motivation, we design a skill-specific evolving planner (SEP) to perform knowledge decoupling, enabling effectively continual learning of novel manipulation skills. Specifically, inspired by [47], we first develop an adaptive language semantic bank to retrieve the skill-specific language semantic embeddings. As for these skill-specific embeddings, SEP can effectively guide the perceiver model $\Theta^{m}_{p}$ to encode multi-modal input from skill-specific latent and low-rank space. It results in learning the novel knowledge through the above skill-shared backbone and a light skill-specific subnetwork.
When observing a novel skill, we utilize a language encoder $\mathcal{E}_{c}$ from CLIP [32] to encode the input language instruction $\mathbf{l}$. Then a skill-specific language semantic embedding $\mathbf{l}_{s}\in\mathbb{R}^{D^{s}}$ can be obtained as: $\mathbf{l}_{s}=\mathcal{E}_{c}(\mathbf{l})$, where $D^{s}$ represents the dimension of $\mathbf{l}_{s}$. To build an adaptive language semantic bank $\mathcal{B}$, we then compensate the semantic embedding $\mathbf{l}_{s}$ via an exponential moving average strategy:
$\mathcal{B}[h,:]=(1-\mathcal{C}_{max})\mathcal{B}[h,:]+\mathcal{C}_{max}\mathbf{l}_{s},$ (8)
where $\mathcal{B}\in\mathbb{R}^{N^{b}\times D^{s}}$ is initialized by $N^{b}$ zero vectors. $\mathcal{C}\in\mathbb{R}^{N^{b}}$ is the cosine similarity matrix between the sentence information $\mathbf{l}_{s}$ and each vector in $\mathcal{B}$. if $\mathcal{C}_{max}>\delta$, we set $h=\mathrm{argmax}(\mathcal{C})$; otherwise, we set $h=\mathrm{nonzero}(\mathcal{B})+1$ and $\mathcal{C}_{max}=1$, where $\delta$ is a fixed threshold, and $\mathrm{nonzero}(\cdot)$ is an operation employed to calculate the number of nonzero vectors in $\mathcal{B}$. In light of this, each skill corresponds to a skill-wise code $h$.
Algorithm 2 Pipeline of Our SEP.
0:  Initialized adaptive language semantic bank $\mathcal{B}$ with $N^{b}$ zero vectors; Initialized dynamic skill-specific latent space $\mathbf{S}=\emptyset$; Initialized low-rank space $\mathbf{W}=\emptyset$; The hyper-parameter $\delta$;
0:  language embedding $\mathbf{l}_{s}$;
0:  $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
1:  Compute cosine similarity matrix $\mathcal{C}$ between $\mathcal{B}$ and $\mathbf{l}_{s}$;
2:  if $\mathcal{C}_{max}>\delta$ then
3:     $h\xleftarrow{}\mathrm{argmax}(\mathcal{C});$
4:     Update $\mathcal{B}$ by Eq. (8);
5:     Return: $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
6:  else
7:     $h\xleftarrow{}\mathrm{nonzero}(\mathcal{B}+1);$
8:     Expand $\mathcal{B}$ by Eq. (8);
9:     Randomly initialize $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
10:     Return: $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
11:  end if
To better learn skill-specific knowledge, we utilize the skill-wise code $h$ to guide the network to perform skill-specific network training. Although existing multi-modal behaviour-cloning methods [39, 53] encode multi-modal input from a skill-shared latent space, we here consider developing a dynamic skill-specific latent space $\mathbf{S}\in\mathbb{R}^{N^{s}\times N^{l}\times D}$ to facilitate skill-specific knowledge learning, where $N^{s}$, $N^{l}$, $D$ denote the number of learned skills, the learnable latent vector quantity and the vector dimension. By following [39], we encode these latent vectors in $\mathbf{S}$ with the multi-modal input to obtain the cross-attention features $\mathbf{F}_{c}\in\mathbb{R}^{(N^{p}+N^{e})\times D}$ as follows:
$\!\!\mathbf{F}_{c}=\rho(\frac{\mathrm{Cat}(\mathbf{p},\mathbf{e})\mathbf{W}_{q}(\mathbf{S}[h,:]\mathbf{W}_{k})^{\top}}{\sqrt{d}})(\mathbf{S}[h,:]\mathbf{W}_{v}),$ (9)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times D}$ are linear projection layers, and $\mathrm{Cat}(\cdot)$ denotes the concatenation operation. $\rho$ indices the softmax function and $d$ is a scaling factor. $\mathbf{p}\in\mathbb{R}^{N^{p}\times D}$ is obtained by concatenating the current state of agent and the input voxelized observation. $\mathbf{e}\in\mathbb{R}^{N^{e}\times D}$ is the encoded embeddings of the input language instruction. To this end, relying on the skill-wise code $h$, NBAgent can continually encode the novel multi-modal input from a skill-specific latent space, which is beneficial to learn skill-specific knowledge from latent space.
Considering the limitation in representing skill-specific knowledge from latent space, we further explore to learn skill-specific knowledge from low-rank space. Specifically, we introduce a low-rank adaptation layer (LoRA) [19] that can learn skill-specific knowledge in an efficient manner. For $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ in each attention block, we design a set of skill-specific LoRA layers to perform skill-specific forward and obtain the final output feature $\mathbf{F}_{x}$ as follows:
$\mathbf{F}_{x}=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{W}_{r}[h,:]=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{W}_{a}[h,:]\mathbf{W}_{b}[h,:],$ (10)
where $\mathbf{W}\in\mathbb{R}^{D\times D}$ represents one of $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ and $\mathbf{X}\in\mathbb{R}^{(N^{p}+N^{e})\times D}$ denotes the input feature for these projection layers. $\mathbf{W}_{r}[h,:]=\mathbf{W}_{a}[h,:]\mathbf{W}_{b}[h,:]$ is a low-rank decomposition, where $\mathbf{W}_{a}\in\mathbb{R}^{D\times N^{r}}$ is initialized in a random Gaussian manner, $\mathbf{W}_{b}\in\mathbb{R}^{N^{r}\times D}$ is initialized by zero and $N^{r}$ is a hyper-parameter controlling the size of LoRA layers. On the one hand, obviously, $\mathbf{W}$ is shared among all skills and expected to learn skill-shared knowledge. On the other hand, each $\mathbf{W}_{r}$ is executed to learn the corresponding skills and perform skill-specific forward with the guidance of skill-wise code $h$, resulting in continually embedding skill-specific knowledge to our NBAgent from low-rank space. Specifically, we summary the process of our SEP in Algorithm 2.
Figure 2: Visualization of prediction results on various manipulation skills between Ours, Ours-w/oSRD, Ours-w/oSEP&SRD, ER and Fine-Tuning.
In conclusion, both SSR and SRD modules above are designed to capture the 3D geometric and semantic information related to the skill, and further transfer skill-shared semantic amongst diverse manipulation skills to tackle past skill forgetting. To perform skill-specific knowledge learning, we develop the SEP to accumulate skill-specific knowledge in latent and low-rank space, thereby effectively learning novel skills. Finally, the objective function of our NBAgent can be simplified as:
$\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{CE}}+\mathcal{L}_{\mathrm{SSR}}+\lambda_{2}\mathcal{L}_{\mathrm{SRD}}.$ (11)
4 Experiments
4.1 Implementation Details
Dataset. By following PerAct [39], we conduct our experiments on RLBench [22] and simulate robot learning in CoppelaSim [34]. To simulate the working scenarios of NBRL, we design two NBRL benchmark datasets, called Kitchen and Living Room. Specifically, Kitchen is constructed by gathering 10 manipulation skills pertinent to kitchen environments, and Living Room consists of 12 manipulation skills associated with living room scenarios. Each manipulation skill includes a training set of 20 episodes and a test set of 25 episodes. Furthermore, these skills involve various variations encompassing randomly sampled attributes such as colors, sizes, counts, placements, and object categories, resulting in a total of 101 distinct variations. In this paper, we use fixed skill ID 1-10 and 1-12 to indicate each skill in different scenarios. The correspondence between each skill ID and the details of skills are provided in Appendix A.1.
Baselines. We conduct the comprehensive evaluation between our NBAgent and the following four methods: PerAct [39], GNFactor [53], Fine-Tuning and ER [7]. PerAct and GNFactor joint train all manipulation skills within one-stage training, for two datasets respectively, referred to as the upperbound. Fine-Tuning achieves continual learning by fine-tuning all parameters in perceiver model on novel skills. ER randomly stores old skills data to memory buff and replay them when detecting novel skills.
Table 1: Comparisons of success rate (%) on Kitchen and Living Room. Red and Blue represents the highest results and runner-up. The number of steps equals the sum of quantity of base skill learning task and continual skill learning tasks.
Comparison Methods 5-5 $(2~{}\mathrm{steps})$ 5-1 $(6~{}\mathrm{steps})$ 6-3 $(3~{}\mathrm{steps})$ 6-2 $(4~{}\mathrm{steps})$
1-5 6-10 All Avg. For. 1-5 6-10 All Avg. For. 1-6 7-12 All Avg. For. 1-6 7-12 All Avg. For.
[…]
Fine-Tuning 15.7 26.4 21.1 38.9 41.1 3.2 20.0 9.6 13.8 39.1 4.4 29.8 17.1 29.0 44.1 6.2 19.1 12.7 24.2 43.9
ER 56.0 25.6 40.8 50.0 3.2 53.6 29.6 41.6 49.7 9.8 43.8 31.1 37.4 42.2 7.7 40.7 30.2 35.4 38.1 17.6
Ours-w/oSEP&SRD 56.0 26.7 41.3 49.1 0.8 56.8 30.4 43.6 50.2 7.1 42.0 33.1 37.6 45.3 12.6 38.4 35.6 37.0 40.9 16.5
Ours-w/oSRD 41.1 38.1 39.6 49.8 18.9 48.0 41.6 44.8 53.9 14.7 41.1 34.4 37.8 45.9 15.6 40.4 39.1 39.8 43.3 19.1
Ours 53.6 36.3 44.9 52.5 6.4 54.4 37.6 46.0 55.2 8.9 44.9 42.2 43.6 47.6 7.7 45.8 35.3 40.6 43.5 10.9
Training Details. In NBRL, following continual learning setting [12], we assume that the agent undergoes initial learning through a set of manipulation skills, referred to as the base skill learning task. Subsequently, skills are added incrementally in steps, characterized as continual skill learning tasks. To valuate the robustness of NBAgent in addressing NBRL problem, we set up four combinations of base skills and new skills: 5-5, 5-1, 6-3, and 6-2, which are highlighted in bold in Tab. 1. For instance, ’5-1’ indicates that the agent is initially trained on a base skill learning task comprising the first five manipulation skills in the skill ID. Subsequently, the agent undergoes five continual skill learning tasks, each involving a new skill, ultimately learning a total of 10 skills in the Kitchen dataset. In addition, the LAMB [51] optimizer is applied for all methods with a initial learning rate of $5.0\times 10^{-4}$ and a batch size of 2. We utilize 100K training iterations for PerAct and GNFactor, 80K for base task training, 20K for incremental task training. We randomly store a fixed 4 episodes of each old skill in $\mathcal{M}$ for ER and Ours. Additional training details are available in Appendix A.2.
Evaluation Metric. Following behaviour-cloning methods [53, 17], we adopt the success score (%) $\mathcal{I}^{m}_{i}$ as the basic indicator for evaluation, where $\mathcal{I}^{m}_{i}$ represents the success score of $i$-th skill tested at $m$-th continual skill learning task. Specifically, after learning the final skill learning task, we compute the mean success score of the skills learned at the base skill learning task (Base) $\mathcal{I}^{M}_{B}$, continual skill learning task (Novel) $\mathcal{I}^{M}_{N}$ and all skill learning tasks (All) $\mathcal{I}^{M}_{A}$. These metrics respectively reflect the robustness of old skill forgetting , the capacity of novel skill learning, as well as its overall performance. Additionally, we introduce a Avg. metric $G$ and For. metric $F$ to measure average performance and skill-wise forgetting rate over the whole NBRL process, where $G=\frac{1}{M}\sum_{m=1}^{M}\mathcal{I}^{m}_{A}$ and $F=\frac{1}{N^{m}}\sum^{N^{m}}_{i=1}\underset{m\in\{1,...,M-1\}}{\mathrm{max}}(\mathcal{I}^{m}_{i}-\mathcal{I}^{M}_{i})$.
Figure 3: Visualization of rendering results in our SSR module. RGB GT denotes the color ground truth, and Semantics GT represents the semantics ground truth extracted by Stable Diffusion model. RGB RD and Semantics RD are the rendering novel view and semantic feature.
4.2 Comparison Performance
We present comparison results between NBAgent and other methods on Kitchen and Living Room datasets in Tabs. 1, 3, 4 and Fig. 2. As shown in Tab. 1, NBAgent significantly outperforms compared methods by $2.5\%\sim 41.4\%$ in terms of Avg metric on four NBRL settings. This indicates that NBAgent effectively addressing old skill forgetting through skill-specific and skill-shared knowledge learning. Furthermore, we presented the skill-wise comparison results in Tabs. 3 and 4. Our model exhibits the highest success rate across all manipulation skills in terms of All. metric, improving by $4.4\%\sim 36.3\%$ and $6.2\%\sim 25.6\%$ on two datasets respectively, compared with Fine-Tuning and ER. This further demonstrates the effectiveness of our model in addressing the NBRL problem. Additionally, our NBAgent is lower than ER by $6.7\%$ in terms of For metrics on Living Room 6-2 setting, which suggests the robust and significant performance of NBAgent over the whole NBRL process. Surprisingly, as shown in Tabs. 3 and 4, our NBAgent performs better than joint training methods, specifically PerAct and GNFactor, providing additional evidence to support the efficacy of our model. The additional comparison results under other settings can be found in Appendix A.3.
Table 2: Comparison results in terms of success rate (%) on Living Room dataset when setting the various size of memory buff $\mathcal{M}$.
Buffer size 6-3 $(2~{}\mathrm{steps})$ 6-2 $(4~{}\mathrm{steps})$
1-6 7-12 All Avg. For. 1-6 7-12 All Avg. For.
$|\mathcal{M}|=2$ 42.0 41.3 41.7 44.9 4.4 32.0 30.7 31.3 37.2 18.8
$|\mathcal{M}|=4$ 44.9 42.2 43.6 47.6 7.7 45.8 35.3 40.6 43.5 10.9
$|\mathcal{M}|=6$ 50.0 36.7 43.3 45.0 2.2 50.0 34.7 42.3 45.4 8.0
4.3 Ablation Studies
Ablation on skill-shared attribute. Compared to Ours, the scores of Ours-w/oSRD on old skills are dropped by $3.8\%\sim 12.5\%$ (i.e., 1-5 and 1-6 in Tab. 1). This indicates that SRD can effectively learn skill-shared knowledge to tackle semantic overlooking on old skills, To evaluate the effectiveness of SSR module, we visualize the rendering results in Fig. 3. It suggests that SSR module can efficiently transfer skill-shared semantics of 3D scenes, thereby achieving an improvement about $0.5\%\sim 3.1\%$ in terms of Avg compared Ours-w/oSEP&SRD with ER in Tab. 1. The above mentioned demonstrates that NBAgent can significantly tackle old skill forgetting by preserving skill-shard semantics of 3D scenes and voxel representation.
Ablation on skill-specific attribute. To evaluate effectiveness of our NBAgent on learning skill-specific knowledge, we eliminate our proposed SEP module and present results in Tabs. 1, 3 and 4. Ours-w/oSRD outperforms Ours-w/oSEP&SRD on novel skills by $1.3\%\sim 11.4\%$, (i.e., 6-10 and 7-12 in Tab. 1), which demonstrates that SEP benefits our model in learning novel skills by performing skill-specific knowledge learning.
Ablation on memory size.We also explore the impact of various sizes of memory buffer $\mathcal{M}$. As shown in Tab. 2, with $|\mathcal{M}|=6$, the forgetting rate of our model notably reduced by $1.1\%\sim 10.8\%$ in comparison to $|\mathcal{M}|=4$ and $2$, respectively. This indicates that increasing the memory size significantly addresses catastrophic forgetting but also incurs a larger memory load.
Table 3: Comparison results on Kitchen under the setting of 5-1.
Skill ID 1 2 3 4 5 6 7 8 9 10 All Imp.
PerAct 96.0 77.3 53.3 30.7 37.3 52.0 2.7 18.7 4.0 73.3 44.5 $\Uparrow$ 1.5
GNFactor 92.0 60.0 70.7 12.0 46.7 52.0 5.3 22.7 10.7 70.7 44.3 $\Uparrow$ 1.7
Fine-Tuning 17.3 0.0 0.0 0.0 0.0 8.0 0.0 0.0 9.3 64.0 9.7 $\Uparrow$ 36.3
ER 90.7 54.7 37.3 54.7 28.0 60.0 5.3 0.0 24.0 62.7 41.6 $\Uparrow$ 4.4
Ours-w/oSEP&SRD 86.7 49.3 34.7 65.3 48.0 66.7 9.3 24.0 5.3 46.7 43.6 $\Uparrow$ 2.4
Ours-w/oSRD 68.0 65.3 50.7 25.3 33.3 76.0 18.7 24.0 22.7 64.0 44.8 $\Uparrow$ 1.2
[…]
Table 4: Comparison results on Living Room under the setting of 6-3.
Skill ID 1 2 3 4 5 6 7 8 9 10 11 12 All Imp.
PerAct 5.3 66.7 0.0 84.0 38.7 13.3 16.0 2.7 41.3 0.0 97.3 6.7 31.0 $\Uparrow$ 12.6
GNFactor 2.7 92.0 1.3 84.0 80.0 28.0 8.0 1.3 46.7 32.0 100 0.0 40.0 $\Uparrow$ 3.6
Fine-Tuning 8.0 18.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 68.0 100 10.7 17.1 $\Uparrow$ 25.6
ER 10.7 58.7 24.0 81.3 76.0 12.0 5.3 4.0 78.7 8.0 89.3 1.3 37.4 $\Uparrow$ 6.2
Ours-w/oSEP&SRD 13.3 74.7 28.0 76.0 52.0 8.0 9.3 12.0 53.3 25.3 96.0 2.7 37.6 $\Uparrow$ 6.0
Ours-w/oSRD 8.0 78.7 6.7 81.3 60.0 12.0 9.3 20.0 52.0 9.3 90.7 25.3 37.8 $\Uparrow$ 5.8
[…]
5 Conclusion
In this paper, we explore a pioneering Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose a novel NBAgent to continually learn skill-wise knowledge. Specifically, we design a skill-shared semantic rendering module and a skill-shared representation distillation module to tackle 3D scene representation overlooking on past encountered skills. Meanwhile, we propose a skill-specific evolving planner to decouple the skill-wise knowledge to effectively learning novel skills and tackling catastrophic forgetting. We develop two NBRL benchmarks to demonstrate the manipulation accuracy, which verifies the effectiveness of our NBAgent against baselines.
Appendix A Supplemental Material
This supplemental material introduces additional details not mentioned in the main paper. Overall, it mainly includes the following aspects:
(a) More Dataset details of Kitchen and Living Room in Sec. A.1.
(b) More implementation details of our proposed NBAgent in Sec. A.1.
(c) More comparison experiments in Sec. A.3.
(d) Analysis of limitations Sec. A.4.
(e) Analysis of societal impact in Sec. A.5.
Table 5: Some important notations and definitions used in this paper.
Notations Definitions Notations Definitions
$\mathcal{T}$ The skill data stream $\mathbf{Y}_{c}$ The input image with ground truth color
$\mathcal{T}^{m}$ The $m$-th continual skill learning task $\mathbf{Y}_{c,t}$ A noisy image
$\mathcal{D}^{m}_{i}$ The demonstration of $i$-th skill in $m$-th task $\mathbf{M}$ The 2D semantic map
$\mathbf{a}$ Action space $\Theta_{u}$ The Stable Diffusion model
$\mathbf{r}$ Structured observation $\mathcal{E}_{v}$ The VAE encoder for image $\mathbf{Y}_{c}$
$\mathbf{l}$ Language instruction $\mathbf{F}_{s}$ The rendered vision-language semantic feature
$\mathbf{Q}^{m}$ The Q-function head in $m$-th continual skill learning task $\hat{\mathbf{F}_{s}}$ The diffusion vision-language semantic feature
$\Theta^{m}_{p}$ The perceiver model in $m$-th continual skill learning task $t$ The diffusion process step
$\Theta^{m}_{n}$ The NeRF model in $m$-th continual skill learning task $\mathbf{l}_{p}$ The language prompt modified from $\mathbf{l}$
$\mathcal{F}_{\Theta^{m}_{n}}$ The neural rending function of $\Theta^{m}_{n}$ $\mathcal{E}_{c}$ The language encoder from CLIP
$\mathbf{v}$ The 3D voxel observation $\mathbf{l}_{s}$ A skill-specific language semantic embedding
$\mathbf{v}_{s}$ The 3D voxel feature $\mathcal{B}$ The adaptive language semantic bank
$\mathbf{x}$ The 3D input point $\mathcal{C}$ The cosine similarity matrix
$\sigma$ Differential density $\mathrm{nonzero}(\cdot)$ An operation employed to calculate the number of nonzero vectors in $\mathcal{B}$
$\mathbf{d}$ The unit viewing direction $h$ A skill-wise code
$\mathbf{r}$ Camera ray $\mathbf{S}$ Dynamic skill-specific latent space
$\mathcal{R}$ The set of all camera rays $\mathbf{F}_{c}$ The cross-attention feature
$\mathbf{o}$ Camera origin $\rho$ The softmax function
$\mathbf{c}$ The color of RGB $\mathbf{e}$ The encoded embeddings of input language instruction
$\mathbf{s}$ The semantic feature $\mathbf{W}$ The MLP projection layer in Transformer block
$\mathbf{C}$ A RGB image $\mathbf{k}^{m}_{j}$ The $j$-th keyframe in a skill, is same with $\mathbf{k}^{m,i}_{j}$
$\hat{\mathbf{C}}$ The pseudo ground truth $\mathbf{v}^{m}_{r,j}$ A keyframe on voxel space get through $\Theta^{m}_{p}$
$N^{d}$ The number of skills in a continual skill learning task $\mathbf{Y}^{m}_{j}$ The ground truth for prediction the next keyframe
$N^{b}$ The number of voctors in $\mathcal{B}$ $\hat{\mathbf{Y}^{m}_{j}}$ The soft label extracted by the perceiver model $\Theta^{m-1}_{p}$
$N^{s}$ The number of learned skills $\mathbf{X}$ The input feature for projection layers in Transformer
$N^{l}$ The learnable latent vector quantity $D$ The dimension of latent vectors in projection layers in Transformer
$N^{p}$ The number of encoded embeddings of input data $D^{s}$ The dimension of $\mathbf{l}_{s}$
$N^{e}$ The number of encoded embeddings of the input language instruction
$N^{r}$ The size of LoRA layers
A.1 Dataset Description
In order to emulate the operational environments of domestic robots, we design two NBRL benchmark datasets using RLBench, called Kitchen and Living Room. In particular, we assemble Kitchen by gathering 10 skills that are pertinent to kitchen settings, and we construct Living Room, which includes 12 skills that are related to living room situations. Each skill includes a training set of 20 episodes and a test set of 25 episodes. We provide examples of brief overviews of these skills and the number of key frames for each skill in Tab. 6 and visualization in Figs. 4-5. In addition, the No. represents the corresponding skill order mentioned in the Experiment section. The Avg. Kf describes the average number of key frames for the skill over 20 episodes in training set.
Table 6: Skill Details in Kitchen & Living Room
No. Skill in Kitchen Language Example Avg. Kf
1 close microwave "close microwave" 2.3
2 meat off grill "take the steak off the grill" 6.0
3 open grill "open the grill" 4.5
4 open wine bottle "open wine bottle" 3.3
5 pick up cup "pick up the red cup" 3.8
6 turn tap "turn left tap" 3.0
7 place wine "stack the wine bottle to the left of the rack" 6.1
8 put knife on "put the knife on the chopping board" 5.1
9 stack wine "stack wine bottle" 7.4
10 take plate off "take plate off the red colored rack" 6.3
No. Skill in Living Room Language Example Avg. Kf
1 close door "close the door" 4.8
2 close laptop lid "close laptop lid" 6.0
3 hang frame "hang frame on hanger" 5.3
4 lamp on "turn on the light" 3.5
5 open drawer "open the bottom drawer" 4.8
6 open window "open left window" 6.0
7 push buttons "push the white button" 4.8
8 put item in drawer "put the item in the top drawer" 14.3
9 put rubbish in bin "put rubbish in bin" 5.0
10 sweep to dustpan "sweep dirt to the tall dustpan" 6.2
11 take usb out "take usb out of computer" 3.3
12 water plants "pour some water on the plant" 6.3
Figure 4: Visualization of manipulation skills in dataset Kitchen.
Figure 5: Visualization of manipulation skills in dataset Living Room.
Table 7: Hyper-parameters used in NBAgent
Variable Name Value
image size 128 × 128 × 3
input voxel size 100 × 100 × 100
batch size $B$ 2
optimizer LAMB
learning rate 0.0005
number of transformer blocks 6
number of sampled points for NBAgent 64
number of latents in Perceiver Transformer 2046
dimension of Stable Diffusion features 512
dimension of CLIP language features 512
hidden dimension of NeRF blocks 512
size of LoRA layers $N^{r}$ 10
base task iterations 80K
incremental task iterations 20K
SSR loss weight $\lambda_{1}$ 0.1
SSR loss function $\mathcal{L}_{SSR}$ MSE-loss
SRD loss weight $\lambda_{2}$ 0.2
SRD loss function $\mathcal{L}_{SRD}$ KL divergence
SRD temperature $\mathcal{T}$ 3
novel skill threshold $\delta$ 0.8
size of memory buffer $\mathcal{M}$ 4
number of base skills in living room 6
number of novel skills in living room 3, 2
number of base skills in kitchen 5
number of novel skills in kitchen 1, 5
Table 8: Comparison results on Kitchen under the setting of 5-5.
Comparison Methods 1 2 3 4 5 6 7 8 9 10 All. Imp.
PerAct 96.0 77.3 53.3 30.7 37.3 52.0 2.7 18.7 4.0 73.3 44.5 $\Uparrow$ 0.4
GNFactor 92.0 60.0 70.7 12.0 46.7 52.0 5.3 22.7 10.7 70.7 44.3 $\Uparrow$ 0.6
Fine-Tuning 10.7 0.0 33.3 33.3 1.3 49.3 6.7 14.7 5.3 56.0 21.1 $\Uparrow$ 23.8
ER 96.0 72.0 56.0 18.7 37.3 45.3 0.0 14.7 1.3 66.7 40.8 $\Uparrow$ 4.1
Ours-w/oSEP&SRD 86.7 74.7 64.0 28.0 26.7 52.0 6.7 17.3 8.0 49.3 41.3 $\Uparrow$ 3.6
Ours-w/oSRD 8.0 81.3 36.0 38.7 41.3 62.7 4.0 46.7 9.3 68.0 39.6 $\Uparrow$ 5.3
[…]
Table 9: Comparison results on Living Room under the setting of 6-2.
Comparison Methods 1 2 3 4 5 6 7 8 9 10 11 12 All. Imp.
PerAct 5.3 66.7 0.0 84.0 38.7 13.3 16.0 2.7 41.3 0.0 97.3 6.7 31.0 $\Uparrow$ 9.6
GNFactor 2.7 92.0 1.3 84.0 80.0 28.0 8.0 1.3 46.7 32.0 100 0.0 40.0 $\Uparrow$ 0.6
Fine-Tuning 0.0 37.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 96.0 18.7 12.7 $\Uparrow$ 27.9
ER 4.0 89.3 17.3 80.0 38.7 14.7 20.0 2.7 60.0 4.0 90.7 4.0 35.4 $\Uparrow$ 5.2
Ours-w/oSEP&SRD 2.7 82.7 14.7 76.0 30.7 24.0 18.7 9.3 76.0 4.0 94.7 10.7 37.0 $\Uparrow$ 3.6
Ours-w/oSRD 12.0 65.3 14.7 82.7 49.3 18.7 28.0 18.7 81.3 6.7 100.0 0.0 39.8 $\Uparrow$ 0.8
[…]
A.2 Implementation Details
As shown in Tab. 7, we give the hyper-parameters used in NBAgent. One agent is trained with two NVIDIA A40 GPU for 2 days (80K + 40K iterations under the 6-3 setting in Living Room) with a batch size of 2. The overall learning rate of the training process is 0.0005, optimized by the LAMB algorithm. All methods employ the same methodology to construct a voxel grid of size $100^{3}$. During the NeRF training, we utilize an additional 19 camera perspectives to furnish supervisory information. For NBRL, we conduct two NBRL settings in each of the two scenarios, which are 5-5, 5-1 in Kitchen, and 6-3, 6-2 in Living Room.
A.3 More Comparison Experiments
Tabs. 8 and 9 show the detailed scores of each model in two scenarios under the NBRL settings of 5-5 and 6-2.With the configuration of memory buffer size $|\mathcal{M}|=4$, the average score of NBAgent across all skills still surpasses other models, approximately between $0.4\%\sim 23.8\%$ and $0.6\%\sim 27.9\%$. This indicates that NBAgent performs the best in solving NBRL problems. Fig. 6 shows the visualization results of more skills among various models. We observe that some skills are not difficult to learn the general actions, but when the skill requires the target object to be in a specific location, the agent has difficulty completing it accurately; for example, ’open the drawer’ is not difficult, but when the requirement is ’open the top drawer’, the performance of NBAgent is better than other models.
Figure 6: Visualization of prediction results on various skills between Ours, Ours-w/oSRD, Ours-w/oSEP&SRD, ER and Fine-Tuing.
A.4 Limitation
While we have found that NBAgent achieves state-of-the-art results on the Never-ending Behavior-cloning Robot Learning (NBRL) problem, we have also identified the following limitations: 1. NBAgent collects keyframes as learning targets using predefined rules in the demonstration dataset, but these rules are not universally applicable to all manipulation tasks. 2.NBAgent cannot continuously learn skills that require continuous actions as it predicts actions based on time-discrete control.
A.5 Social Impact
In practical multi-task robot learning, we explore the Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose a novel NBAgent. The goal is to continually learn new skills in an incremental manner, allowing for unique customized operational abilities. For example, robots can learn new skills by imitating specific demonstration datasets tailored to the owner’s needs. Additionally, we address catastrophic forgetting in continual skill learning by transferring skill-shared semantic of 3D scenes using a skill-shared semantic rendering module and a skill-shared representation distillation module.
Typically, robots need strong learning capabilities to adapt to various work scenarios effectively. Having the ability for continual learning allows robots to provide better user experiences. Furthermore, this approach enables robots to learn specific operational skills based on the diverse needs of different user groups. For instance, household robots need to learn operational tasks in various contexts. Therefore, the proposed Never-ending Behavior-cloning Robot Learning problem is worth investigating.